OpenStudy (anonymous):

Rahim purchased twenty dozen of toys at the rate of Rs 375 per dozen. He sold each one of them at the rate of Rs 33. What was his percentage profit?

OpenStudy (anonymous):

What's the equation you would write to calculate his costs? Twenty dozen toys at a cost of 375 per dozen.

OpenStudy (anonymous):

even i dont know!!

OpenStudy (anonymous):

If one dozen costs 375, how much does 20 dozen cost?

OpenStudy (anonymous):

375*20

OpenStudy (ajprincess):

ya right.

OpenStudy (anonymous):

Cool. So then how much does he sell them for? 20 dozen at a price of 33 each.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Once you get how much he sold them for, subtract out what he paid for them (the first number you calculated). Then take that number and divide it by what he paid for them. That's his profit percentage.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

total no. of toys=20*12=240 CP=20*375=7500 SELLING PRICE=240*33=7920 PRFIT=7920-7500=420 PRFIT%=420/7500*100=5.6%

OpenStudy (anonymous):

@aroo1993 CHECK MY ANSWER